OpenStudy (anonymous):

Fan and Medal A boat traveled 12 miles with the current and 15 miles against it. The boat's average speed is 10 miles per hour. A. Let c represent the current. Write a rational expression for the time with the current. B. Let c represent the current. Write a rational expression for the time against the current. C. Write an equation to represent the total time. D. Simplify the equation written in Part C. E. Find the total time if the speed of the current is 3 mph. Make sure to label your answer.

OpenStudy (anonymous):

@epoweritheta

OpenStudy (perl):

I have seen this question ago

OpenStudy (neer2890):

\[average speed=\frac{ total . distance. covered }{ total. time . taken }\] \[10=\frac{ 12+15 }{ \frac{ 12 }{ x+c } +\frac{ 15 }{ x-c }}\] where x is the speed of boat and c is the speed of current.

OpenStudy (neer2890):

so \[t1=\frac{ 12 }{ x+c }\] is the time with the current. \[t2=\frac{ 15 }{ x-c }\] is the time against the current and t1+t2 is the total time taken.

OpenStudy (neer2890):

now put the value of c=3 in the equation of average speed and solve for x. find t1 and t2 for total time.

OpenStudy (perl):

any questions